4a-3[2y+6(a-3y)-5a]
Answer by tommyt3rd(5050)

4a-3[2y+6(a-3y)-5a]=
4a-3[2y+6a-18y-5a]=
4a-3[a-16y]=
4a-3a+48y=
a+48y